# Occupancy-SLAM: An Efficient and Robust Algorithm for Simultaneously Optimizing Robot Poses and Occupancy Map

### v2

@@ II-C Joint Optimization of Poses and Maps
State vector of optimization, $\mathbf{x}={[{\mathbf{x}^{P}}^{\top},{\mathbf{x}^{M}}^{\top}]}^{\top}$. $\mathbf{x}^{l}$ and $\mathbf{x}^{s}$
@@ 12                  Evaluate $F(\mathbf{x})$ at $\mathbf{x}^{P}(k)$ and $\mathbf{x}^{M}(k)$
13                  Solve $\mathbf{J}^{\top}\mathbf{W}\mathbf{J}\mathbf{\Delta}(k)=-\mathbf{J}^{\top}\mathbf{W}F(\mathbf{x})$, where $\mathbf{\Delta}(k)={[{\mathbf{\Delta}^{P}(k)}^{\top},{\mathbf{\Delta}^{M}(k)}^{\top}]}^{\top}$
@@ V-B Our Multi-resolution Joint Optimization Strategy
Initialized by the odometry inputs or scan matching, we perform Algorithm 1 to quickly obtain relatively accurate poses. The state vector in the first stage is ${\mathbf{x}}^{l}={[{\mathbf{x}^{P}}^{\top},{{\mathbf{x}^{lM}}}^{\top}]}^{\top}$, where $\mathbf{x}^{lM}$ includes all occupancy values at the cell vertices of the low-resolution map $\mathbb{M}^{l}$. In this stage, the hit map, observation information, coefficient matrix, weight matrix, and resolution are represented as $\mathbb{N}^{l},\mathbb{S}^{l},\mathbf{A}^{l}$, $\mathbf{W}^{l}$, and $s^{l}$, respectively.
@@ 6      Calculate the convoluted map $\mathbb{C}$ with kernel size $q$
In the second stage, the state vector is represented as $\mathbf{x}^{s}={[{{\mathbf{x}}^{P}}^{\top},{\mathbf{x}^{sM})}^{\top}]}^{\top}$ where $\mathbf{x}^{sM}$ includes all occupancy values at cell vertices of the selected high-resolution map $\mathbb{M}^{s}$. We perform Algorithm 3 using poses obtained from the first stage as initial guesses and observations $\mathbb{S}^{s}$ to refine poses. The NLLS optimization problem in the second stage can be formulated similarly as (17). Additionally, the differences in the Jacobian calculation between Algorithm 1 and Algorithm 3 are described in Appendix -E.
@@ VI Submap Joining
In Section V, we introduced a multi-resolution joint optimization strategy to efficiently solve our NLLS problem. For large-scale occupancy SLAM with long robot trajectories, the number of poses to optimize can be very large. To make the computational complexity dependent only on the environment size rather than the trajectory length, in this section we propose an occupancy submap joining method. The key idea is to reduce the number of poses that need to be optimized to the number of local submaps.
@@ VI-A Inputs and Outputs of Submap Joining Problem
We first separate the observation information into multiple parts and perform Algorithm 4 to build several submaps. The inputs of submap joining problem are a sequence of local occupancy submaps. Let us denote the $n_{L}+1$ submaps as $\mathbb{M}_{L}=\{\mathbb{M}_{L_{0}},\cdots,\mathbb{M}_{L_{n_{L}}}\}$ and the associated coordinate frames of these local occupancy maps are denoted as $\{\mathbf{x}^{P}_{0},\cdots,\mathbf{x}^{P}_{n_{L}}\}$, where $\mathbb{M}_{L_{i_{L}}}$ and $\mathbf{x}^{P}_{i_{L}}$represents the $i_{L}$th local occupancy map and its associated coordinate frame. In addition, the global occupancy map is represented as $\mathbb{M}_{G}=\{M_{G}(\mathbf{m}^{G}_{00}),\cdots,{M}_{G}\left(\mathbf{m}^{G}_{c_{w}^{G}c_{h}^{G}}\right)\}$. Both the global map and local maps follow the same definition as described in Section III-B. The outputs of submap joining problem are the optimal solution of the local submap coordinate frames and the optimal global occupancy map.
@@ VI-B NLLS Formulation of Submap Joining Problem
$\mathbf{x}_{G}=[{\mathbf{x}^{P}_{L}}^{\top},{{\mathbf{x}^{M}_{G}}}^{\top}]^{\top},$ (24)
@@
$g(\mathbf{x}_{G})=\sum\limits_{i_{L}=0}^{n}\sum\limits_{wh\in\mathbb{S}^{L}_{i_{L}}}\left\|\omega(i_{L},\mathbf{m}^{G}_{wh}){M}_{G}(\mathbf{m}^{G}_{wh})-{M}_{L_{i_{L}}}(\mathbf{p}_{i_{L}}^{wh})\right\|^{2},$ (26)
where $\mathbb{S}^{L}_{i_{L}}$ represents the set of indices of cell vertices in the global occupancy map $\mathbb{M}_{G}$ that are projected onto the local submap $\mathbb{M}_{L_{i_{L}}}$.
@@ VII-C Assessment of Robustness to Initial Guess
Moreover, for all practical datasets, we additionally add random zero-mean uniform distribution noises ([2[-2[ - 2 m, $2$ m]]]] for translation and [0.5[-0.5[ - 0.5 rad, $0.5$ rad]]]] for rotation) to the poses obtained from Cartographer as the initial guess. The initial occupancy maps obtained by using the noisy initial poses are shown in Fig. 13(a). Fig. 13(b) shows the remapped occupancy grid maps using our optimized poses, and Fig. 13(c) shows the point cloud maps using our optimized poses. This experiment shows that our approach can converge from initial guesses with significant errors and also generate good results.
@@ VII-D Discussion about the Effectiveness of Different Stages
One potential question is whether using only Algorithm 1 with a high-resolution map would yield even better results. To investigate this, we compared our full approach with Algorithm 1 using a high-resolution map. We tested three initialization: (1) Algorithm 1 (High, O/S): initialization using odometry inputs or scan matching; (2) Algorithm 1 (High, Carto): initialization using Cartographer’s poses (as proposed in our conference paper [10]); and (3) Algorithm 1 (High, First): initialization using the poses obtained by our first stage.
@@
It is also worth noting that our full method utilizes the selected high-resolution map for optimization in the second stage. As shown in Fig. 14, in certain experiments, the accuracy of our full approach surpasses that of the optimization using the full high-resolution map (Algorithm 1 (High, First)). A possible explanation is that once relatively accurate poses and occupancy maps are obtained, the dropped cell vertices and corresponding observations contain little information. Retaining all cells and corresponding observations for optimization may prevent the algorithm from finding the optimal solution, as all are observation items given uniform weights. Another reason could be the smoothing term in (10), which spreads occupancy values to unknown cell vertices, potentially introducing errors that affect the convergence of the optimization.
@@
The reduction in time consumption stems from both the multi-resolution strategy, which reduces time per iteration, and the fewer iterations needed in the second stage due to the selected high-resolution map. Our experiments show that only about two iterations are needed in the second stage with the selected high-resolution map, fewer than in Algorithm 1 (High, First). This is likely because the selected high-resolution map focuses on critical states, with observations containing the most relevant information, enabling faster convergence.
@@ VII-E Ablation Study on the Resolution Ratio
We assess accuracy and computational time using three simulation datasets, with the resolution in the second stage fixed at $s^{h}=0.05$ m. The resolution ratios $r$ between the first and second stages are set to 2, 5, 10, and 20, respectively. To ensure consistency, a fixed selection range of $d=1.5$ m is applied uniformly across all datasets.
@@ VII-G Computational Complexity Analysis
The Gauss-Newton method for solving the joint optimization of local maps and poses in (17) and submap joining in (26) primarily depends on calculating Jacobian $\mathbf{J}$, and constructing and solving the sparse linear system in (18) [38]. We analyze each part’s complexity separately due to differences in the NLLS formulation.
@@
For the local map and poses joint optimization problem, the objective function consists of the observation term, the odometry term, and the smoothing term. Let ${\lambda(\mathbb{S})}$ denotes the number of sampling points $\mathbb{S}$, then the number of items in the objective function is $\mathfrak{d}_{row}=\lambda(\mathbb{S})+3(n-1)+2{c_{w}}{c_{h}}+c_{w}+c_{h}$, and the state vector size is $\mathfrak{d}_{col}=3n+(c_{w}+1)(c_{h}+1)$. Considering Jacobian of the smoothing term $\mathbf{J}_{S}$ can be pre-calculated before optimization, the number of non-zero elements of Jacobian matrix that need to be computed for each iteration is $\mathfrak{d}_{J}=7\lambda(\mathbb{S})+6(n-1)$. Therefore, for each iteration, the computation complexity of Jacobian calculation, constructing (18) and solving (18) is $\mathcal{O}(\mathfrak{d}_{J})$, $\mathcal{O}(\mathfrak{d}_{J}\mathfrak{d}_{col})$, and $\mathcal{O}({\mathfrak{d}^{3}_{col}})$, respectively. Therefore, the total computation complexity per iteration for the local map and poses joint optimization problem is $\mathcal{O}(\mathfrak{d}_{J}+\mathfrak{d}_{J}{\mathfrak{d}_{col}}+\mathfrak{d}^{3}_{col})$. Due to our proposed multi-resolution joint optimization strategy and keyframe selection, both $\mathfrak{d}_{J}$ and $\mathfrak{d}_{col}$ remain small during the first and second stages of optimization, making the computation time for this part manageable.
@@
For our submap joining algorithm, the number of observations depends on the total number of cell vertices of the global occupancy map observed in each submap, denoted $\mathfrak{d}_{obs}^{G}$. Considering that some cell vertices will be observed repeatedly under different submaps, this number slightly exceeds the number of non-unknown cell vertices in the global map. Thus, the number of non-zero elements of Jacobian matrix is $\mathfrak{d}_{J}^{G}=4\mathfrak{d}_{obs}^{G}$, and the state vector size is $\mathfrak{d}_{col}^{G}=3n_{L}+(c_{w}^{G}+1)(c_{h}^{G}+1)$. The computation complexity per iteration is $\mathcal{O}(\mathfrak{d}_{J}^{G}+\mathfrak{d}_{J}^{G}{\mathfrak{d}_{col}^{G}}+{\mathfrak{d}_{col}^{G}}^{3})$. Although the global occupancy map tends to be relatively large, the sub-matrix of Hessian w.r.t. the global occupancy map is diagonal. To speed up computation, we apply the Schur complement [57] to make the normal equation solving highly efficient.
@@
Finally, we evaluate the time consumption of our method using both all frames and selected keyframes in large-scale environments to support our computation complexity analysis and compare it to Cartographer. For Museum b2 dataset, Cartographer takes 1424 seconds, while our method takes 1250 seconds when using all frames and 363 seconds with selected key frames. For C3 dataset, Cartographer takes 610 seconds, while our method takes 742 seconds with all frames and 236 seconds with selected key frames. In our total time consumption, the submap joining method consumes less than 10 seconds on both datasets. It can be seen that the time consumption of our method is comparable to that of Cartographer when all frames are used and much lower than that of Cartographer when selected key frames are used. These results demonstrate the efficiency of our multi-resolution joint optimization strategy and submap joining approach.
@@ VIII-A Extension of the Algorithms to 3D Case
The submap joining problem in 3D remains largely similar to the 2D case, except that the projection relation extends from 2D-2D to 3D-3D, enabling the solution of 6 DoF poses and the 3D global occupancy map in the NLLS problem (26).
@@ VIII-B1 Evaluation metrics and state-of-the-art methods
We evaluate our method’s performance in 3D using absolute trajectory error for poses, aligning and comparing results with ground truth via EVO [58], as used in [49, 22]. In all the experiments, we use the odometry information provided by the dataset as initialization if it is available. Otherwise, we use FAST-LIO2 [59] to obtain the odometry information. To evaluate our method, we compare our method against state-of-the-art methods: BALM2 [48], HBA [49], and Voxgraph [33]. BALM2 optimizes the planar feature parameters of the point cloud and the robot’s poses. HBA proposes a hierarchical bundle adjustment to optimize the consistency of the planar surfaces of point clouds and robot poses. Voxgraph builds SDF-based submaps from point clouds, uses SDF-to-SDF registration for relative submap measurements, and incrementally optimizes submap frames. HBA and Voxgraph can deal with large-scale environments, while BALM2 focuses on normal-scale environments.

### v3

@@ II-C Joint Optimization of Poses and Maps
State vector of optimization, $\mathbf{x}={\left[{\mathbf{x}^{P}}^{\top},{\mathbf{x}^{M}}^{\top}\right]}^{\top}$. $\mathbf{x}^{l}$ and $\mathbf{x}^{s}$
@@ 12                  Evaluate $F(\mathbf{x})$ at $\mathbf{x}^{P}(k)$ and $\mathbf{x}^{M}(k)$
13                  Solve $\mathbf{J}^{\top}\mathbf{W}\mathbf{J}\mathbf{\Delta}(k)=-\mathbf{J}^{\top}\mathbf{W}F(\mathbf{x})$, where $\mathbf{\Delta}(k)={\left[{\mathbf{\Delta}^{P}(k)}^{\top},{\mathbf{\Delta}^{M}(k)}^{\top}\right]}^{\top}$
@@ V-B Our Multi-resolution Joint Optimization Strategy
Initialized by the odometry inputs or scan matching, we perform Algorithm 1 to quickly obtain relatively accurate poses. The state vector in the first stage is ${\mathbf{x}}^{l}={\left[{\mathbf{x}^{P}}^{\top},{{\mathbf{x}^{lM}}}^{\top}\right]}^{\top}$, where $\mathbf{x}^{lM}$ includes all occupancy values at the cell vertices of the low-resolution map $\mathbb{M}^{l}$. In this stage, the hit map, observation information, coefficient matrix, weight matrix, and resolution are represented as $\mathbb{N}^{l},\mathbb{S}^{l},\mathbf{A}^{l}$, $\mathbf{W}^{l}$, and $s^{l}$, respectively.
@@ 6      Calculate the convoluted map $\mathbb{C}$ with kernel size $q$
In the second stage, the state vector is represented as $\mathbf{x}^{s}={\left[{{\mathbf{x}}^{P}}^{\top},{\mathbf{x}^{sM})}^{\top}\right]}^{\top}$ where $\mathbf{x}^{sM}$ includes all occupancy values at cell vertices of the selected high-resolution map $\mathbb{M}^{s}$. We perform Algorithm 3 using poses obtained from the first stage as initial guesses and observations $\mathbb{S}^{s}$ to refine poses. The NLLS optimization problem in the second stage can be formulated similarly as (17). Additionally, the differences in the Jacobian calculation between Algorithm 1 and Algorithm 3 are described in Appendix -E.
@@ VI Submap Joining
In Section V, we introduced a multi-resolution joint optimization strategy to efficiently solve our NLLS problem. For large-scale occupancy SLAM with long robot trajectories, the number of poses to optimize can be very large. To make the computational complexity mainly depend on the environment size rather than the trajectory length, in this section we propose an occupancy submap joining method. The key idea is to reduce the number of poses that need to be optimized to the number of local submaps.
@@ VI-A Inputs and Outputs of Submap Joining Problem
We first separate the observation information into multiple parts and perform Algorithm 4 to build several submaps. The inputs of submap joining problem are a sequence of local occupancy submaps. Let us denote the $n_{L}+1$ submaps as $\mathbb{M}_{L}=\left\{\mathbb{M}_{L_{0}},\cdots,\mathbb{M}_{L_{n_{L}}}\right\}$ and the associated coordinate frames of these local occupancy maps are denoted as $\{\mathbf{x}^{P}_{0},\cdots,\mathbf{x}^{P}_{n_{L}}\}$, where $\mathbb{M}_{L_{i_{L}}}$ and $\mathbf{x}^{P}_{i_{L}}$represents the $i_{L}$th local occupancy map and its associated coordinate frame. In addition, the global occupancy map is represented as $\mathbb{M}_{G}=\left\{M_{G}(\mathbf{m}^{G}_{00}),\cdots,{M}_{G}\left(\mathbf{m}^{G}_{c_{w}^{G}c_{h}^{G}}\right)\right\}$. Both the global map and local maps follow the same definition as described in Section III-B. The outputs of submap joining problem are the optimal solution of the local submap coordinate frames and the optimal global occupancy map.
@@ VI-B NLLS Formulation of Submap Joining Problem
$\mathbf{x}_{G}=\left[{\mathbf{x}^{P}_{L}}^{\top},{{\mathbf{x}^{M}_{G}}}^{\top}\right]^{\top},$ (24)
@@
$g(\mathbf{x}_{G})=\sum\limits_{i_{L}=0}^{n}\sum\limits_{wh\in\mathbb{S}^{G}_{i_{L}}}\left\|\omega(i_{L},\mathbf{m}^{G}_{wh}){M}_{G}(\mathbf{m}^{G}_{wh})-{M}_{L_{i_{L}}}(\mathbf{p}_{i_{L}}^{wh})\right\|^{2},$ (26)
where $\mathbb{S}^{G}_{i_{L}}$ represents the set of indices of cell vertices in the global occupancy map $\mathbb{M}_{G}$ that are projected onto the local submap $\mathbb{M}_{L_{i_{L}}}$.
@@ VII-C Assessment of Robustness to Initial Guess
Moreover, for all practical datasets, we additionally add random zero-mean uniform distribution noises ([2[-2[ - 2 m, $2$ m]]]] for translation and [0.5[-0.5[ - 0.5 rad, $0.5$ rad]]]] for rotation) to the poses obtained from Cartographer as the initial guess. The initial occupancy maps obtained by using the noisy initial poses are shown in Fig. 13(a). Fig. 13(b) shows the remapped occupancy grid maps using our optimized poses, and Fig. 13(c) shows the point cloud maps using our optimized poses. This experiment shows that our approach can converge from very poor initial guesses and generate good results.
@@ VII-D Discussion about the Effectiveness of Different Stages
One question is whether using only Algorithm 1 with a high-resolution map would yield even better results. To investigate this, we compared our full approach with Algorithm 1 using a high-resolution map. We tested three initialization: (1) Algorithm 1 (High, O/S): initialization using odometry inputs or scan matching; (2) Algorithm 1 (High, Carto): initialization using Cartographer’s poses (as proposed in our conference paper [10]); and (3) Algorithm 1 (High, First): initialization using the poses obtained by our first stage.
@@
It is also worth noting that our full method utilizes the selected high-resolution map for optimization in the second stage. As shown in Fig. 14, in certain experiments, the accuracy of our full approach surpasses that of the optimization using the full high-resolution map (Algorithm 1 (High, First)). A possible explanation is that once relatively accurate poses and occupancy maps are obtained, the dropped cell vertices and corresponding observations contain little information. Retaining all cell vertices and corresponding observations for optimization may prevent the algorithm from finding the optimal solution, as all observation items are given uniform weights. Another reason could be the smoothing term in (10), which spreads occupancy values to unknown cell vertices, potentially introducing errors that affect the convergence of the algorithm.
@@
The reduction in time consumption stems from both the multi-resolution strategy, which reduces time per iteration, and the fewer iterations needed in the second stage due to the selected high-resolution map. Our experiments show that only about two iterations are needed in the second stage with the selected high-resolution map, fewer than in Algorithm 1 (High, First). This is probably because the selected high-resolution map focuses on critical states, with observations containing the most relevant information, enabling faster convergence.
@@ VII-E Ablation Study on the Resolution Ratio
We assess the accuracy and computational time using three simulation datasets, with the resolution in the second stage fixed at $s^{h}=0.05$ m. The resolution ratios $r$ between the first and second stages are set to 2, 5, 10, and 20, respectively. To ensure consistency, a fixed selection range of $d=1.5$ m is applied uniformly across all datasets.
@@ VII-G Computational Complexity Analysis
The computational cost of the Gauss–Newton method for solving the joint optimization of local maps and poses in (17), as well as submap joining in (26), is primarily determined by computing the Jacobian matrix $\mathbf{J}$, constructing and solving the sparse linear system in (18) [38]. Since the formulations of the NLLS problems differ between these two methods, we analyze their computational complexities separately.
@@
In the case where odometry inputs are not considered, the objective function of the joint optimization problem within each local submap $\mathbb{M}_{L_{i_{L}}}$ consists of the observation and smoothing terms. Let ${\lambda(\mathbb{S}^{L}_{i_{L}})}$ and $n^{L}_{i_{L}}$ denote the number of sampling points $\mathbb{S}^{L}_{i_{L}}$ and the number of poses within the $i_{L}$-th submap, respectively. Let $c^{{i_{L}}}_{w}\times c^{{i_{L}}}_{h}$ represent the size of the $i_{L}$-th submap. Then, the number of residual terms in the objective function is $\mathfrak{d}_{row}=\lambda(\mathbb{S}^{L}_{i_{L}})+3(n^{L}_{i_{L}}-1)+2c^{{i_{L}}}_{w}c^{{i_{L}}}_{h}+c^{{i_{L}}}_{w}+c^{{i_{L}}}_{h}$, and the size of state vector is $\mathfrak{d}_{col}=3n^{L}_{i_{L}}+(c^{{i_{L}}}_{w}+1)(c^{{i_{L}}}_{h}+1)$. Considering Jacobian of the smoothing term $\mathbf{J}_{S}$ can be pre-calculated, the number of non-zero elements of the Jacobian matrix to be computed per iteration is $\mathfrak{d}_{J}=7\lambda(\mathbb{S}^{L}_{i_{L}})+6(n^{L}_{i_{L}}-1)$. Therefore, the per-iteration computation complexities of the Jacobian calculation, constructing (18) and solving (18) are $\mathcal{O}(\mathfrak{d}_{J})$, $\mathcal{O}(\mathfrak{d}_{J}\mathfrak{d}_{col})$, and $\mathcal{O}({\mathfrak{d}^{3}_{col}})$, respectively. The total computation complexity per iteration for the optimization problem is $\mathcal{O}(\mathfrak{d}_{J}+\mathfrak{d}_{J}{\mathfrak{d}_{col}}+\mathfrak{d}^{3}_{col})$. Due to our proposed multi-resolution joint optimization strategy and keyframe selection, both $\mathfrak{d}_{J}$ and $\mathfrak{d}_{col}$ remain small during the first and second stages of optimization, making the computation time manageable.
@@
For our submap joining algorithm, the number of observations depends on the total number of cell vertices in the global occupancy map that are observed across all submaps, denoted as $\mathfrak{d}_{obs}^{G}$. Considering that some cell vertices are observed repeatedly from different submaps, this number slightly exceeds the total number of non-unknown cell vertices in the global map. The number of non-zero elements in Jacobian matrix is therefore $\mathfrak{d}_{J}^{G}=4\mathfrak{d}_{obs}^{G}$, and the size of state vector is $\mathfrak{d}_{col}^{G}=3n_{L}+(c_{w}^{G}+1)(c_{h}^{G}+1)$. The computation complexity per iteration is $\mathcal{O}(\mathfrak{d}_{J}^{G}+\mathfrak{d}_{J}^{G}{\mathfrak{d}_{col}^{G}}+{\mathfrak{d}_{col}^{G}}^{3})$. Although the global occupancy map is typically large, the sub-matrix of Hessian w.r.t. the global occupancy map is diagonal. To speed up computation, we apply the Schur complement [57] to make the normal equation solving highly efficient.
@@
To validate our computational complexity analysis, we evaluate the time consumption of our method using both all frames and selected key frames in large-scale environments and compare it with Cartographer. On Museum b2 dataset, Cartographer takes 1424 seconds, while our method takes 1250 seconds when using all frames and 363 seconds with selected key frames. On C3 dataset, Cartographer requires 610 seconds, while our method takes 742 seconds with all frames and 236 seconds with selected key frames. Notably, in both datasets, the submap joining step in our method accounts for less than 10 seconds of the total runtime. These results demonstrate that our method achieves runtime comparable to Cartographer when using all frames, and significantly lower runtime when using selected key frames. This confirms the effectiveness and efficiency of our multi-resolution joint optimization strategy and submap joining approach.
@@ VIII-A Extension of the Algorithms to 3D Case
The submap joining problem in 3D remains largely similar to the 2D case, with the primary difference being that the projection relation extends from 2D-2D to 3D-3D, enabling the estimation of 6 DoF poses and the 3D global occupancy map in the NLLS problem (26).
@@ VIII-B1 Evaluation metrics and state-of-the-art methods
We evaluate our method in 3D using absolute trajectory error (MAE and RMSE), aligning estimated trajectories with ground truth via EVO [58], as used in [49, 22]. For all experiments, we use the dataset-provided odometry as initialization when available. Otherwise, we use FAST-LIO2 [59] to obtain the odometry. Our method is compared against state-of-the-art methods, including BALM2 [48], HBA [49], and Voxgraph [33]. BALM2 jointly optimizes the planar feature parameters extracted from point clouds and the robot’s poses. Building on this idea, HBA introduces a hierarchical bundle adjustment framework to improve scalability for large-scale environments. Voxgraph builds SDF-based submaps from point clouds, performs SDF-to-SDF registration for relative alignment, and incrementally optimizes submap frames. HBA and Voxgraph target large-scale environments, while BALM2 is designed for normal-scale environments.